# Fork-Merge Decoding: Enhancing Multimodal Understanding in Audio-Visual Large Language Models

### v1

@@ Abstract
The goal of this work is to enhance balanced multimodal understanding in audio-visual large language models (AV-LLMs) by addressing modality bias without requiring additional training. In current AV-LLMs, audio and video features are typically processed jointly in the decoder. While this strategy facilitates unified multimodal understanding, it may introduce modality bias, where the model tends to over-rely on one modality due to imbalanced training signals. To mitigate this, we propose Fork-Merge Decoding (FMD), a simple yet effective inference-time strategy that requires no additional training or architectural modifications. FMD first performs modality-specific reasoning by processing audio-only and video-only inputs through the early decoder layers (a fork phase), and then merges the resulting hidden states for joint reasoning in the remaining layers (a merge phase). This approach promotes balanced modality contributions and leverages complementary information across modalities. We evaluate our method on two representative AV-LLMs, VideoLLaMA2 and video-SALMONN, using three benchmark datasets. Experimental results demonstrate consistent performance improvements on tasks focused on audio, video, and combined audio-visual reasoning, demonstrating the effectiveness of inference-time interventions for robust multimodal understanding.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated superior performance in various text-centric tasks such as problem solving, translation, and summarization achiam2023gpt ; brown2020language ; liu2023summary ; thoppilan2022lamda ; wei2021finetuned ; wei2022chain ; zhao2023survey . Building on these successes in text processing, LLMs have evolved to handle additional modalities, including images alayrac2022flamingo ; chen2023shikra ; dai2023instructblip ; huang2023language ; li2023blip ; iu2023visual ; yu2024rlhf ; zhang2023llama ; zhu2023minigpt , videos lin2023video ; maaz2023video , and audio huang2024audiogpt ; rubenstein2023audiopalm ; tang2023salmonn , giving rise to multimodal LLMs (MLLMs). These models process diverse modalities through separate encoders and integrate their outputs within a language model decoder, achieving remarkable performance across a wide range of tasks. Among these, audio-visual LLMs (AV-LLMs) are particularly notable for their ability to jointly integrate visual and auditory information, enabling more sophisticated reasoning and closer alignment with human multimodal understanding cheng2024videollama ; sun2024video ; zhang2023video .
@@
To effectively leverage pretrained LLM decoders in AV-LLMs, various fusion strategies have been proposed to integrate audio and visual information. One common approach chen2023vast ; cheng2024videollama ; chowdhury2024meerkat ; han2024onellm ; lyu2023macaw ; panagopoulou2023x ; ye2024cat ; zhan2024anygpt ; zhang2023video ; zhao2023chatbridge is token-wise fusion, where audio and visual features are extracted by separate encoders and then concatenated along the sequence dimension before being fed into the decoder as a continuous input sequence. Several studies chowdhury2024meerkat ; ye2024cat additionally introduce adapter modules that facilitate interaction between audio and visual features before they are passed into the LLM. Another approach han2023imagebind ; su2023pandagpt ; sun2024video is channel-wise fusion, in which modality-specific features are concatenated along the channel dimension to form a unified representation. In most current AV-LLM architectures, the decoder receives both audio and visual inputs simultaneously, which raises a potential concern: if the model finds one modality easier to interpret—perhaps due to better alignment with its pretraining objectives—it may over-rely on that modality, resulting in modality-specific hallucinations leng2024curse .
@@
Figure 1: Attention weight analysis in VideoLLaMA2 cheng2024videollama on AVHBench dataset sung2024avhbench . We analyze 100 samples and examine the attention weights from the last decoder layer, focusing on the last token in the question. The attention is disproportionately assigned to video inputs over audio, indicating a modality bias. Our proposed FMD method mitigates this gap by promoting more balanced contributions from both modalities.
@@
To investigate this possibility, we begin by analyzing the attention weight distributions over audio-visual inputs using 100 samples from the AVHBench dataset sung2024avhbench . Our analysis focuses on the final decoder layer of VideoLLaMA2 cheng2024videollama , examining the attention weights of the last token to quantify the relative attention allocated to each modality. As shown in Figure 1, the vanilla decoding setup—representing the model’s default inference behavior—exhibits a clear bias toward video inputs, with attention disproportionately concentrated on visual features over audio. This observation aligns with findings from recent studies guan2024hallusionbench ; leng2024curse ; nishimura2024audio ; wang2024mitigating , which report that LLM-based decoders in multimodal settings often exhibit modality bias. These studies highlight that such imbalances can lead to hallucinations or flawed reasoning, emphasizing the need to mitigate modality bias for more robust and balanced multimodal understanding.
@@
To address this issue, we propose Fork-Merge Decoding (FMD), which is a simple yet effective strategy that enhances multimodal understanding without altering the AV-LLM architecture or requiring additional training. The core idea of FMD is to divide the decoding phase into two stages: a fork phase and a merge phase, designed to improve both unimodal and multimodal understanding. In the fork phase, the original multimodal input is split into two unimodal branches by zeroing out either the visual or auditory modality while retaining the text question. Each branch is independently processed through the early layers of the pretrained AV-LLM, yielding modality-specific hidden representations. In the merge phase, these representations are combined and passed through the remaining decoder layers to enable joint reasoning. This separation enables the model to first attend to unimodal cues in isolation before integrating them for complementary multimodal understanding. As shown in Figure 1, FMD effectively reduces modality bias and balances audio-visual contributions. Furthermore, since recent AV-LLMs commonly adopt either token-wise or channel-wise concatenation for audio-visual fusion, we propose a generalized decoding strategy that is compatible with both. This unified approach improves performance across a variety of models, regardless of their fusion mechanism. Recent AV-LLM architectures and their fusion methods are summarized in Supp. A.
@@
Furthermore, we evaluate the effectiveness of FMD by applying it to two recent AV-LLMs: VideoLLaMA2, which performs token-wise fusion, and video-SALMONN sun2024video , which adopts channel-wise fusion. With FMD, both models achieve improved decoding performance across three widely used audio-visual benchmarks: AVQA yang2022avqa , MUSIC-AVQA li2022learning , and AVHBench sung2024avhbench . Notably, FMD enhances performance not only in tasks that emphasize a single modality but also in tasks that require balanced reasoning across both modalities. These results demonstrate that FMD effectively supports multimodal understanding by fully leveraging the information from each modality during inference.
@@ 2 Preliminaries
Input processing. In many existing AV-LLMs chen2023vast ; cheng2024videollama ; chowdhury2024meerkat ; lyu2023macaw ; panagopoulou2023x ; sun2024video ; ye2024cat ; zhan2024anygpt ; zhang2023video ; zhao2023chatbridge , audio and video inputs are first processed independently through their respective encoders before being jointly fed into the LLM decoder. Visual inputs are typically handled by a vision encoder that processes each sampled frame to generate frame-level embeddings capturing spatial information. Audio inputs are transformed by an audio encoder into semantic representations that encode acoustic and prosodic features of the signal. Textual inputs, such as instructions or questions, are tokenized and embedded using the tokenizer and embedding layer associated with the base language model.
For token-wise fusion-based AV-LLMs, video inputs are first transformed into a fixed set of $M$ embeddings, denoted as $\mathbf{x}^{v}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}\}$. Similarly, audio inputs are encoded into $N$ embeddings, represented as $\mathbf{x}^{a}=\{\mathbf{x}_{M+1},\ldots,\mathbf{x}_{M+N}\}$. These visual and audio embeddings are concatenated along the token dimension to form an initial multimodal sequence. The question text is tokenized and embedded into $L$ text embeddings using the language model’s text embedding layer, yielding $\mathbf{x}^{l}=\{\mathbf{x}_{M+N+1},\ldots,\mathbf{x}_{M+N+L}\}$. The final input sequence $\mathbf{x}$ is constructed by concatenating the visual, audio, and text embeddings: $\mathbf{x}=\mathbf{x}^{v}\oplus\mathbf{x}^{a}\oplus\mathbf{x}^{l}$, where $\oplus$ denotes token-wise concatenation. This produces a sequence of length $M+N+L$. For simplicity, we omit instruction tokens, which are typically prepended to the input sequence.
@@
In contrast, channel-wise fusion-based AV-LLMs first project both visual and audio features into embeddings of a fixed length $U$, and then concatenate them along the channel (feature) dimension to form joint audio-visual embeddings: $\mathbf{x}_{i}^{av}=[\mathbf{x}_{i}^{v};\mathbf{x}_{i}^{a}]$, where $\mathbf{x}_{i}^{v}$ and $\mathbf{x}_{i}^{a}$ denote the visual and audio embeddings, respectively. These audio-visual embeddings are subsequently combined with the $L$ text embeddings through token-wise concatenation. The resulting input sequence is defined as $\mathbf{x}=\{\mathbf{x}_{1}^{av},\ldots,\mathbf{x}_{U}^{av},\mathbf{x}_{U+1}^{l},\ldots,\mathbf{x}_{U+L}^{l}\}$, which is then fed into the decoder for further processing.
Figure 2: Overview of the Fork-Merge Decoding (FMD) pipeline. The AV-LLM takes video frames, an audio waveform, and a question prompt as input. In the fork phase, FMD masks either the video or audio modality while preserving the question prompt, enabling independent reasoning over each modality. After $L_{\text{fork}}$ decoder layers in $\phi$, the merge phase combines the resulting hidden states, $\mathbf{h}_{\text{fork}}^{v}$ and $\mathbf{h}_{\text{fork}}^{a}$, using an importance weight $\alpha$ derived from attention distributions over unmasked modality tokens; the complementary weight $(1-\alpha)$ is applied to masked counterparts. This decoding scheme enables the pretrained AV-LLM to integrate both audio and visual cues effectively, allowing it to resolve questions that may be ambiguous or misleading when relying on a single modality.
@@
Decoding. Both token-wise and channel-wise fusion-based AV-LLMs generate outputs from the input sequence $\mathbf{x}$ using an autoregressive decoding strategy, where each token is predicted by attending to previously generated tokens under a causal mask. At timestep $t$, the model generates the next token $\mathbf{y}_{t}$ conditioned on the input sequence $\mathbf{x}$ and the previously generated tokens $\mathbf{y}_{<t}$:
$\mathbf{y}_{t}$ $\sim p(\mathbf{y}_{t}|\mathbf{x},\mathbf{y}_{<t})\ \propto\exp\left(\text{logit}(\mathbf{y}_{t}|\mathbf{x},\mathbf{y}_{<t})\right).$ (1)
At each decoding step, the model computes hidden states conditioned on the causal context $\mathbf{y}_{<t}$. These hidden states are then passed through a linear projection layer, commonly known as the vocabulary head, to produce logits over the vocabulary. A softmax function is applied to the logits to produce a probability distribution over candidates for the next token.
3 Enhancing Audio-Visual understanding with Fork-Merge Decoding
@@
This section introduces Fork-Merge Decoding (FMD), a general decoding framework applicable to two representative fusion architectures in AV-LLMs: token-wise fusion (Section 3.1) and channel-wise fusion (Section 3.2). An overview of the process is illustrated in Figure 2.
@@ 3.1 Decoding with Token-wise Fusion in AV-LLMs
Masking strategy for input sequences. In token-wise fusion models like VideoLLaMA2 cheng2024videollama , the input sequence is composed of $\mathbf{x}=\mathbf{x}^{v}\oplus\mathbf{x}^{a}\oplus\mathbf{x}^{l}$, where $\mathbf{x}^{v}$, $\mathbf{x}^{a}$, and $\mathbf{x}^{l}$ denote visual, audio, and language embeddings, respectively. To enable modality specific processing for audio and video while preserving the textual question, we construct two masked variants of the input:
$\mathbf{x}^{v\text{[masked]}}={\mathbf{x}}^{\neg v}\oplus\mathbf{x}^{a}\oplus\mathbf{x}^{l},\quad\mathbf{x}^{a\text{[masked]}}=\mathbf{x}^{v}\oplus{\mathbf{x}}^{\neg a}\oplus\mathbf{x}^{l},$ (2)
where ${\mathbf{x}}^{\neg v}$ and ${\mathbf{x}}^{\neg a}$ denote the modality-masked embeddings for vision and audio, respectively. These are obtained by zeroing out the corresponding video frames or audio waveforms at the input level. This preserves original embedding shapes and positions while removing content information.
@@
$\mathbf{h}^{v}_{\text{fork}}=\phi_{\leq L_{\text{fork}}}(\mathbf{x}^{v\text{[masked]}}),\quad\mathbf{h}^{a}_{\text{fork}}=\phi_{\leq L_{\text{fork}}}(\mathbf{x}^{a\text{[masked]}}),$ (3)
@@
where $\mathbf{h}^{v}_{\text{fork}}$ and $\mathbf{h}^{a}_{\text{fork}}$ denote the intermediate hidden states obtained from the vision-masked and audio-masked inputs, respectively. This design ensures that the model does not observe both audio and visual inputs simultaneously in the early stages, allowing it to focus on how each modality individually relates to the textual prompt. By reasoning over each modality in isolation, the model is less likely to become biased toward the more dominant or easier-to-interpret modality.
@@
Merge processing. After the $L_{\text{fork}}$ layers, the hidden states $\mathbf{h}^{v}_{\text{fork}}$ and $\mathbf{h}^{a}_{\text{fork}}$ are fused and passed through the remaining transformer layers. Each hidden state has a sequence length of $M+N+L$, corresponding to the visual, audio, and text tokens. The fusion is performed by summing the corresponding embeddings at each modality position, and is formally defined as follows:
$\mathbf{h}_{\text{merge}}[\mathcal{V}]$ $=(1-\alpha)\cdot\mathbf{h}^{v}_{\text{fork}}[\mathcal{V}]+\alpha\cdot\mathbf{h}^{a}_{\text{fork}}[\mathcal{V}],$ (4)
$\mathbf{h}_{\text{merge}}[\mathcal{A}]$ $=\alpha\cdot\mathbf{h}^{v}_{\text{fork}}[\mathcal{A}]+(1-\alpha)\cdot\mathbf{h}^{a}_{\text{fork}}[\mathcal{A}],$
$\mathbf{h}_{\text{merge}}[\mathcal{L}]$ $=\tfrac{1}{2}\left(\mathbf{h}^{v}_{\text{fork}}[\mathcal{L}]+\mathbf{h}^{a}_{\text{fork}}[\mathcal{L}]\right),$
@@
where $\mathcal{V}=[1{:}M]$, $\mathcal{A}=[M{+}1{:}M{+}N]$, and $\mathcal{L}=[M{+}N{+}1{:}M{+}N{+}L]$ denote the index ranges corresponding to the visual, audio, and language embeddings, respectively. The final merged representation $\mathbf{h}_{\text{merge}}$ is then constructed by concatenating the modality-specific segments along the token dimension as $\mathbf{h}_{\text{merge}}=\mathbf{h}_{\text{merge}}[\mathcal{V}]\oplus\mathbf{h}_{\text{merge}}[\mathcal{A}]\oplus\mathbf{h}_{\text{merge}}[\mathcal{L}]$. Here, $\alpha$ serves as a modality-specific weighting factor that controls the relative contribution of each unmasked modality in the fusion. We refer to this approach as attention-guided fusion, which is described in detail in the following section.
@@
Attention-guided fusion. To determine the fusion weight $\alpha$, we leverage the attention matrix $\mathbf{A}^{\text{final}}\in\mathbb{R}^{T\times T}$ from the final transformer layer. Specifically, we focus on the attention vector of the last token, $\mathbf{a}_{\text{last}}=\mathbf{A}^{\text{final}}_{T,:}\in\mathbb{R}^{T}$, which plays a critical role in next-token prediction. Using masked inputs $\mathbf{x}^{v[\text{masked}]}$ and $\mathbf{x}^{a[\text{masked}]}$, we compute the attention-based contribution of each unmasked modality by comparing the attention mass it receives relative to the masked version as follows:
$\alpha_{a}=\frac{\sum_{i\in\mathcal{A}}\mathbf{a}_{\text{last}}^{v\text{[masked]}}[i]}{\sum_{i\in\{\mathcal{V}\cup\mathcal{A}\}}\mathbf{a}_{\text{last}}^{v\text{[masked]}}[i]},\quad\alpha_{v}=\frac{\sum_{i\in\mathcal{V}}\mathbf{a}_{\text{last}}^{a\text{[masked]}}[i]}{\sum_{i\in\{\mathcal{V}\cup\mathcal{A}\}}\mathbf{a}_{\text{last}}^{a\text{[masked]}}[i]},$ (5)
where $\mathbf{a}_{\text{last}}^{v\text{[masked]}}$ and $\mathbf{a}_{\text{last}}^{a\text{[masked]}}$ represent the attention weight distributions of the final token when the visual or audio modality is masked, respectively. Finally, we compute the average attention mass assigned to the unmasked modality in each of the video-masked and audio-masked branches:
$\alpha=\frac{1}{2}(\alpha_{a}+\alpha_{v}).$ (6)
The aggregated weight $\alpha$ is used to interpolate between $\mathbf{h}^{v}_{\text{fork}}$ and $\mathbf{h}^{a}_{\text{fork}}$ in Eq. (4). In practice, instead of computing $\alpha$ for each input, we estimate representative values by sampling 100 samples from the AVHBench sung2024avhbench dataset. These fixed weights are then used consistently across all experiments.
@@
Decoding. The merged hidden state is then forwarded through the remaining transformer layers:
$\mathbf{h}_{\text{final}}=\phi_{>L_{\text{fork}}}(\mathbf{h}_{\text{merge}}),$ (7)
producing the final prediction logits. By delaying modality fusion to deeper layers—where individual representations become semantically richer through the fork phase—our method enhances multimodal understanding while mitigating issues caused by modality imbalance.

### v2

@@ Abstract
The goal of this work is to enhance balanced multimodal understanding in audio-visual large language models (AV-LLMs) by addressing modality bias without additional training. In current AV-LLMs, audio and video features are typically processed jointly in the decoder. While this strategy facilitates unified multimodal understanding, it may introduce modality bias, where the model tends to over-rely on one modality due to imbalanced training signals. To mitigate this, we propose Fork-Merge Decoding (FMD), a simple yet effective inference-time strategy that requires no additional training or architectural modifications. FMD first performs modality-specific reasoning by processing audio-only and video-only inputs through the early decoder layers (fork), and then merges the resulting hidden states for joint reasoning in the remaining layers (merge). This separation allows each modality to be emphasized in the early stages while encouraging balanced contributions during integration. We validate our method on three representative AV-LLMs—VideoLLaMA2, video-SALMONN, and Qwen2.5-Omni—using three benchmark datasets. Experimental results show consistent gains in audio, video, and audio-visual reasoning tasks, highlighting the effectiveness of inference-time interventions for robust and efficient multimodal understanding.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated superior performance in various text-centric tasks such as problem solving, translation, and summarization (Achiam et al., 2023; Brown et al., 2020; Liu et al., 2023b; Thoppilan et al., 2022; Wei et al., 2022a; b; Zhao et al., 2023a). Building on these successes in text processing, LLMs have evolved to handle additional modalities, including images (Alayrac et al., 2022; Chen et al., 2023a; Dai et al., 2023; Huang et al., 2023; Li et al., 2023a; Liu et al., 2023a; Yu et al., 2024; Zhang et al., 2024; Zhu et al., 2023b), videos (Lin et al., 2024; Maaz et al., 2024), and audio (Huang et al., 2024; Rubenstein et al., 2023; Tang et al., 2024), giving rise to multimodal LLMs (MLLMs). These models process diverse modalities through separate encoders and integrate their outputs within a decoder language model, achieving remarkable performance across a wide range of tasks. Among these, audio-visual LLMs (AV-LLMs) are particularly notable for their ability to jointly integrate visual and auditory information, supporting more sophisticated reasoning and achieving closer alignment with human multimodal perception (Cheng et al., 2024; Sun et al., 2024; Zhang et al., 2023; Xu et al., 2025).
@@
To effectively leverage pretrained LLM decoders in AV-LLMs, various fusion strategies have been proposed to integrate audio and visual information. One common approach (Chen et al., 2023b; Cheng et al., 2024; Chowdhury et al., 2024; Han et al., 2024; Lyu et al., 2023; Panagopoulou et al., 2023; Ye et al., 2024; Zhan et al., 2024; Zhang et al., 2023; Zhao et al., 2023b; Xu et al., 2025) is token-wise fusion, where audio and visual features are extracted by separate encoders and then concatenated along the sequence dimension before being fed into the decoder as a continuous input sequence. Several studies (Chowdhury et al., 2024; Ye et al., 2024) additionally introduce adapter modules that facilitate interaction between audio and visual features before they are passed into the LLM. Another approach (Han et al., 2023; Su et al., 2023; Sun et al., 2024) is channel-wise fusion, in which modality-specific features are concatenated along the channel dimension to form a unified representation. In most current AV-LLM architectures, the decoder receives both audio and visual inputs simultaneously, which raises a potential concern: if the model finds one modality easier to interpret—perhaps due to better alignment with its pretraining objectives—it may over-rely on that modality, leading to modality bias and modality-specific hallucinations (Leng et al., 2024a).
@@
Figure 1: Attention weight analysis in VideoLLaMA2 on the AVHBench dataset. We analyze 100 samples and examine the attention weights from the last decoder layer, focusing on the final token of the question. Attention is disproportionately allocated to video inputs over audio, revealing a modality bias. Our proposed FMD method reduces this gap by encouraging more balanced contributions from both modalities.
@@
To investigate this possibility, we begin by analyzing the attention weight distributions over audio-visual inputs using 100 samples from the AVHBench dataset (Sung-Bin et al., 2025). Our analysis examines the final decoder layer of VideoLLaMA2 (Cheng et al., 2024), focusing on the attention weights of the last token. Since the token is critical for predicting the next token, we use it to quantify the relative attention allocated to each modality. As shown in Figure 1, the vanilla decoding setup, which reflects the default inference behavior of the model, exhibits a clear bias toward video inputs, with attention disproportionately concentrated on visual features over audio. This observation aligns with findings from recent studies (Guan et al., 2024; Leng et al., 2024a; Nishimura et al., 2024; Wang et al., 2024), which report that MLLM decoders often exhibit modality bias. These studies highlight that such imbalances can lead to hallucinations or flawed reasoning, emphasizing the need to mitigate modality bias for more balanced multimodal understanding.
@@
To address this issue, we propose Fork-Merge Decoding (FMD), which is a simple yet effective strategy that enhances multimodal understanding without altering the AV-LLM architecture or requiring additional training. The core idea of FMD is to divide the decoding phase into two stages: a fork phase and a merge phase, designed to improve both unimodal and multimodal understanding. In the fork phase, the original multimodal input is split into two unimodal branches by zeroing out either the visual or auditory modality while retaining the text question. Each branch is processed independently through the initial layers of the pretrained AV-LLM, producing modality-specific hidden representations without requiring additional full forward passes. In the merge phase, these representations are combined and passed through the remaining decoder layers. This separation enables the model to first attend to unimodal cues in isolation before integrating them for complementary multimodal understanding. As shown in Figure 1, FMD reduces the attention weight on video inputs by 14% and increase the weight on audio inputs by 7%. This adjustment balances the modality bias while still preserving the attributes of the pretrained model. Building on this, since recent AV-LLMs commonly adopt either token-wise or channel-wise concatenation for audio-visual fusion, we propose a generalized decoding strategy that is compatible with both fusion methods. This unified approach improves performance across a variety of models, regardless of their fusion mechanism.
@@
Furthermore, we evaluate the effectiveness of FMD by applying it to three recent AV-LLMs: VideoLLaMA2 (Cheng et al., 2024) and Qwen2.5-Omni (Xu et al., 2025) (token-wise fusion) and video-SALMONN (Sun et al., 2024) (channel-wise fusion). Applying FMD leads to consistent performance improvements in all baselines across three widely used audio-visual benchmarks: AVQA (Yang et al., 2022), MUSIC-AVQA (Li et al., 2022), and AVHBench (Sung-Bin et al., 2025). Notably, FMD enhances performance not only in tasks that emphasize a single modality but also in tasks that require balanced reasoning across both modalities. These results show that FMD enhances multimodal understanding by fully utilizing information from each modality during inference.
@@ 2 Preliminaries
Input processing. Most existing AV-LLMs (Chen et al., 2023b; Cheng et al., 2024; Chowdhury et al., 2024; Lyu et al., 2023; Panagopoulou et al., 2023; Sun et al., 2024; Ye et al., 2024; Zhan et al., 2024; Zhang et al., 2023; Zhao et al., 2023b) process audio and video separately before feeding them into an LLM decoder. Visual inputs are encoded frame by frame into spatial embeddings, while audio signals are mapped to semantic representations capturing acoustic and prosodic cues. Textual instructions or questions are tokenized and embedded by a tokenizer of language model.
@@
For token-wise fusion models, video frames are transformed into $M$ embeddings ${\bm{x}}^{v}=\{{\bm{x}}_{1},\ldots,{\bm{x}}_{M}\}$ and audio into $N$ embeddings ${\bm{x}}^{a}=\{{\bm{x}}_{M+1},\ldots,{\bm{x}}_{M+N}\}$. These are concatenated with $L$ text embeddings ${\bm{x}}^{l}=\{{\bm{x}}_{M+N+1},\ldots,{\bm{x}}_{M+N+L}\}$ to form the final sequence ${\bm{x}}={\bm{x}}^{v}\oplus{\bm{x}}^{a}\oplus{\bm{x}}^{l}$ of length $M+N+L$, omitting instruction tokens for simplicity. For channel-wise fusion models, visual and audio features are projected to a fixed length $U$ embeddings and then concatenated along the channel dimension to form joint audio-visual embeddings, ${\bm{x}}_{i}^{av}=[{\bm{x}}_{i}^{v};{\bm{x}}_{i}^{a}]$. The resulting set ${\bm{x}}_{i}^{av}=\{{\bm{x}}_{1}^{av},\ldots,{\bm{x}}_{U}^{av}\}$ is combined with $L$ text embeddings via token-wise concatenation, ${\bm{x}}=\{{\bm{x}}_{1}^{av},\ldots,{\bm{x}}_{U}^{av},{\bm{x}}_{U+1}^{l},\ldots,{\bm{x}}_{U+L}^{l}\}$, which is then fed into the decoder.
@@
Decoding. Both token-wise and channel-wise fusion-based AV-LLMs generate outputs from the input sequence $\bm{x}$} using an autoregressive decoding strategy, where each token is predicted by attending to previously generated tokens under a causal mask. At each decoding step $t$, the model generates the next token ${\bm{y}}_{t}$ conditioned on the input sequence $\bm{x}$}, which includes video, audio, and text prompts, as well as the previously generated tokens ${\bm{y}}_{<t}$, with ${\bm{y}}_{t}\sim p({\bm{y}}_{t}|{\bm{x}},{\bm{y}}_{<t})\ \propto\exp\left(\text{logit}({\bm{y}}_{t}|{\bm{x}},{\bm{y}}_{<t})\right).$
Figure 2: Overview of the Fork-Merge Decoding pipeline. The AV-LLM takes video frames, an audio waveform, and a question prompt as input. In the fork phase, FMD masks one modality while preserving the question, enabling independent reasoning. After $L_{\text{fork}}$ decoder layers, the merge phase combines ${\bm{h}}_{\text{fork}}^{\neg v}$ and ${\bm{h}}_{\text{fork}}^{\neg a}$ with an attention-derived weight $\alpha$, and the merged representation is processed by the remaining layers to generate answers with balanced multimodal understanding.
3 Fork-Merge Decoding for Audio-Visual Understanding
@@
This section introduces Fork-Merge Decoding (FMD) for token-wise (Section 3.1) and channel-wise (Section 3.2) fusion in AV-LLMs. An overview of the process is illustrated in Figure 2.
@@ 3.1 Decoding with token-wise fusion in AV-LLMs
Input masking strategy. In token-wise fusion models like VideoLLaMA2 (Cheng et al., 2024), the input sequence is composed of ${\bm{x}}={\bm{x}}^{v}\oplus{\bm{x}}^{a}\oplus{\bm{x}}^{l}$, where ${\bm{x}}^{v}$, ${\bm{x}}^{a}$, and ${\bm{x}}^{l}$ denote visual, audio, and language embeddings, respectively. To enable modality-specific processing for audio and video while preserving the textual question, we create two masked input variants as follows:
${\bm{x}}^{v\text{[masked]}}={{\bm{x}}}^{\neg v}\oplus{\bm{x}}^{a}\oplus{\bm{x}}^{l},\quad{\bm{x}}^{a\text{[masked]}}={\bm{x}}^{v}\oplus{{\bm{x}}}^{\neg a}\oplus{\bm{x}}^{l},$ (1)
where ${{\bm{x}}}^{\neg v}$ and ${{\bm{x}}}^{\neg a}$ denote the modality-masked embeddings for vision and audio, respectively. These are obtained by zeroing out the corresponding video frames or audio waveforms at the input level. This preserves original embedding shapes and positions while removing content information.
@@
${\bm{h}}^{\neg v}_{\text{fork}}=\phi_{\leq L_{\text{fork}}}({\bm{x}}^{v\text{[masked]}}),\quad{\bm{h}}^{\neg a}_{\text{fork}}=\phi_{\leq L_{\text{fork}}}({\bm{x}}^{a\text{[masked]}}),$ (2)
@@
where ${\bm{h}}^{\neg v}_{\text{fork}}$ and ${\bm{h}}^{\neg a}_{\text{fork}}$ denote the intermediate hidden states obtained from the vision-masked and audio-masked inputs, respectively. This design ensures that the model does not observe both audio and visual inputs simultaneously in the early stages, allowing it to focus on how each modality individually relates to the textual prompt. By reasoning over each modality in isolation, the model is less likely to become biased toward the more dominant or easier-to-interpret modality.
@@
Merge processing. After the $L_{\text{fork}}$ layers, the hidden states ${\bm{h}}^{\neg v}_{\text{fork}}$ and ${\bm{h}}^{\neg a}_{\text{fork}}$ are fused and passed through the remaining transformer layers. Each hidden state has a sequence length of $M+N+L$, corresponding to the visual, audio, and text tokens. The fusion is performed by summing the corresponding embeddings at each modality position, and is formally defined as follows:
${\bm{h}}_{\text{merge}}[\mathcal{V}]$ $=(1-\alpha)\cdot{\bm{h}}^{\neg v}_{\text{fork}}[\mathcal{V}]+\alpha\cdot{\bm{h}}^{\neg a}_{\text{fork}}[\mathcal{V}],$ (3)
${\bm{h}}_{\text{merge}}[\mathcal{A}]$ $=\alpha\cdot{\bm{h}}^{\neg v}_{\text{fork}}[\mathcal{A}]+(1-\alpha)\cdot{\bm{h}}^{\neg a}_{\text{fork}}[\mathcal{A}],$
${\bm{h}}_{\text{merge}}[\mathcal{L}]$ $=\tfrac{1}{2}\left({\bm{h}}^{\neg v}_{\text{fork}}[\mathcal{L}]+{\bm{h}}^{\neg a}_{\text{fork}}[\mathcal{L}]\right),$
@@
where $\mathcal{V}=[1{:}M]$, $\mathcal{A}=[M{+}1{:}M{+}N]$, and $\mathcal{L}=[M{+}N{+}1{:}M{+}N{+}L]$ denote the index ranges corresponding to the visual, audio, and language embeddings, respectively. The final merged representation ${\bm{h}}_{\text{merge}}$ is then constructed by concatenating the modality-specific segments along the token dimension as ${\bm{h}}_{\text{merge}}={\bm{h}}_{\text{merge}}[\mathcal{V}]\oplus{\bm{h}}_{\text{merge}}[\mathcal{A}]\oplus{\bm{h}}_{\text{merge}}[\mathcal{L}]$. Here, $\alpha$ is a fusion weight for unmasked segments, reflecting their relative contribution when combining masked and unmasked segments. We refer to this approach as attention-guided fusion and describe in detail in the following section.
@@
Attention-guided fusion. To determine the fusion weight $\alpha$, we use the attention matrix $\mathbf{A}^{\text{final}}\in\mathbb{R}^{T\times T}$ from the final transformer layer. Specifically, we focus on the attention vector of the last token, $\mathbf{a}_{\text{last}}=\mathbf{A}^{\text{final}}_{T,:}\in\mathbb{R}^{T}$, which is critical in next-token prediction and follows the approach in prior studies (Huo et al., 2025; Song et al., 2024). By leveraging the two masked branches ${\bm{h}}^{\neg v}_{\text{fork}}$ and ${\bm{h}}^{\neg a}_{\text{fork}}$, we compute the attention-based contributions of unmasked segments by summing the attention mass over the corresponding regions relative to the total mass, which are then used as $\alpha$:
$\alpha=\frac{\sum_{i\in\mathcal{A}}\mathbf{a}_{\text{last}}^{v[\text{masked}]}[i]+\sum_{i\in\mathcal{V}}\mathbf{a}_{\text{last}}^{a[\text{masked}]}[i]}{\sum_{i\in\mathcal{V}\cup\mathcal{A}}(\mathbf{a}_{\text{last}}^{v[\text{masked}]}[i]+\mathbf{a}_{\text{last}}^{a[\text{masked}]}[i])},$ (4)
where $\mathbf{a}_{\text{last}}^{v\text{[masked]}}$ and $\mathbf{a}_{\text{last}}^{a\text{[masked]}}$ represent the attention weight distributions of the final token from ${\bm{h}}^{\neg v}_{\text{fork}}$ and ${\bm{h}}^{\neg a}_{\text{fork}}$, respectively. The final attention weight $\alpha$ is then obtained as the fraction of unmasked attention over the total attention in Eq. 4 and is used to interpolate between ${\bm{h}}^{\neg v}_{\text{fork}}$ and ${\bm{h}}^{\neg a}_{\text{fork}}$ in Eq. 3. In practice, instead of computing $\alpha$ for each data point, we estimate a representative value by sampling 100 random examples from the AVHBench (Sung-Bin et al., 2025) dataset. This approach is motivated by two considerations: (1) computing a separate $\alpha$ for each sample increases inference time, as it requires two additional full forward passes to obtain $\mathbf{a}_{\text{last}}^{v\text{[masked]}}$ and $\mathbf{a}_{\text{last}}^{a\text{[masked]}}$, (2) noisy sample-specific $\alpha$ values can act as outliers, causing performance drops, as shown by the results labeled Attention (Adaptive) in Table 3. The resulting representative $\alpha$ is then applied consistently across all experiments, including datasets beyond AVHBench, to verify its generalizability.
@@
Decoding. The merged hidden state is then forwarded through the remaining transformer layers: ${\bm{h}}_{\text{final}}=\phi_{>L_{\text{fork}}}({\bm{h}}_{\text{merge}}),$ producing the final prediction logits. By delaying modality fusion to deeper layers, where individual representations become semantically richer through the fork phase, our method enhances multimodal understanding while mitigating issues caused by modality imbalance.